OpenStudy (anonymous):

I really need help with this question!! Please!!! Find four consecutive even integers, such that the product of the first and second integers minus 8 is equal to the product of the third and fourth integers

OpenStudy (goformit100):

Sir/Ma'am, A Warm Welcome To Open Study. Sir/Ma'am please show your Working for the question you have posted.

OpenStudy (aravindg):

x,x+2,x+4,x+6 x(x+2)-8=(x+4)(x+6) Does that look helpful? Can you do it now?

OpenStudy (anonymous):

I already have that step I am not sure what my next step would be... The only difference is I made my variable "n". Can you please help me a little further. I would really appreciate it.

OpenStudy (aravindg):

Open up the brackets and solve the quadratic for x.

OpenStudy (anonymous):

okay i got n^2+2n-8=n^2+6n+4n+24 n^2+2n-8=n^2+10n+24 -8n-32=0 -8(n+4)=0 n=-4 Therefore n=-4 n+2=> -4+2=-2 n+4=> -4+4=0 n+6=> -4+6=2 is this correct??

OpenStudy (goformit100):

▬Sir/Ma'am, A Warm Welcome To Open Study. Sir/Ma'am please show your Working for the question you have posted.▬

OpenStudy (aravindg):

Yes it is correct :)

OpenStudy (anonymous):

Thank you!! :D